CYANOPOLYYNE CATIONS AS CARRIERS OF A SET OF DIFFUSE INTERSTELLAR BANDS

\underline{JAMES K. G. WATSON}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario, Canada K1A 0R6.

It is proposed that a set of diffuse interstellar bands (DIBs) can be attributed to the B̃^{2}Π − X̃^{2}Π electronic transitions of the cyanopolyyne cations, HC_{2n+1}N^{+}, based on comparisons with laboratory observations in Ne matrices by Forney et al.^{1} For HC_{7}N^{+} the DIB λ6614 is assigned as the 0_{0}^{0} origin band, and λλ6426, 6196, 5982, and 5850 to transitions to upper-state fundamental vibrations. The three-peaked structure of λ6614 under high resolution^{2,3} is assigned as P and R branches of the two subbands ^{2}Π_{1/2} −^{2}Π_{1/2} and ^{2}Π_{3/2} −^{2}Π_{3/2}, with the central pair of branches overlapping. The 0_{0}^{0} band of HC_{9}N^{+} is assigned to λ7562, with vibrational bands at λλ7358, 6919, and 6521, while the 0_{0}^{0} band of HC_{11}N^{+} is assigned to λ8531, with vibrational bands too weak to identify. These are all relatively sharp DIBs, with full widths at half maximum (FWHMs) in the range 1.7–3.1 cm^{−1}. Assignments are less satisfactory for shorter members of the series. The 0_{0}^{0} bands of HC_{5}N^{+} and HC_{3}N^{+} may be λλ5797 and 5110, respectively, but the former seems to have a rather narrow P R separation under high resolution, although its FWHM of 2.9 cm^{−1} lies in the above range, while the latter is much broader (FWHM = 45 cm^{−1}) than higher members of the series. The first member of the series, HCN^{+}, does not have a … π^{3}π^{4} ← … π^{4}π^{3} electronic transition of this type.

1. D. Forney, P. Freivogel, J. Fulara, and J. P. Maier, J. Chem. Phys., 102, 1510 (1995).

2. P. J. Sarre, J. R. Miles, T. H. Kerr, R. E. Hibbins, S. J. Fossey, and W. B. Somerville, Mon. Not. Roy. Astron. Soc. 277, L41 (1995).

3. P. Ehrenfreund and B. H. Foing, Astron. Astrophys., 307, L25 (1996).